OpenStudy (anonymous):

Omar is 66 years old and Emily is 20 years old. How many years will it take until Omar is only 3 times as old as Emily?

OpenStudy (anonymous):

3 20+3=23 66+3=69 23*3=69

OpenStudy (mathstudent55):

This problem can be solved with an equation or by trial and error. 20 is just a little less than a third of 66, so it won't be many years until Omar is only 3 times as old as Emily. Just write a table. Start with their ages now. In each line, add one year to their ages until Omar's age is 3 times Emily's age. I started the table for you below. Just keep adding lines.